Use ration exponents to write x^1/5*y^1/2*z^1/8 as a single radical expression
Answer by Theo(13342)

the least common multiple for the denominator of all the exponents is 40.
1/5 = 8/40
1/2 = 20/40
1/8 = 5/40
use that to combine into one radical as follows:
x^(1/5) * y^(1/2) * z^(1/8) is equal to:
x^(8/40) * y^(20/40) * z^(5/40) which is equal to:
(x^8 * y^20 * z^5) ^ (1/40).
in radical form this is equal to:
the 40th root of (x^8 * y^20 * z^5).